Please solve these two problems without using EXCEL and provide your work. Write down the formulas you used to solve the problems and solve the problems.

1. Assume your company purchased $145,305 worth of office equipment (7-year useful life) and sold it for $32,999 at the end of year 7. What is the book value of the property at the end of year 2 using MACRS depreciation and a 5% MARR. (ANSWER IS 88,956)

2. The Steinbrenner family donated $3.7 million for the construction of a new stadium on the MIT campus in 1955. At the same time the family donated an additional $3.7 million as a gift to be used for the construction of a future stadium. If MIT invested the gift in an account with an interest rate of 9.7%, what was the value of the gift in millions of actual dollars in 2010 assuming that the average inflation rate was 4.7%? Note please express your answer in millions of dollars rather than dollars (e.g., 1.2 not 1,200,000) (ANSWER IS 601.96)
Answer #1
Answer 1)
In MACRS salvage value is not needed
Using MACRS tables for depreciation
useful life 7 years
Year 1 depreciation
145305 14.30% 20778.62 Book value at year end 1= 145305-20778.62 124526.4
Year 2 depreciation
145305 24.50% 35599.73 Book value at year end 1= 124526.4-35599.73 88926.67
Answer 2)
FV in 1955 for 2010 will be
=3.7 * (1.097)55 601.96